A 12% acid sol. Was mixd wth a 16% acid sol. To make 40L of a 15% acid sol. How much of the 12% and 16% acid solutions were used?
Answer by Maths68(1474)

Solution A
Amount = x
Concentration =12% =0.12
Solution B
Amount = 40-x
Concentration =16% = 0.16
Resultant Solution
Amount =40
Concentration =15%=0.15
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
(x)(0.12)+(40-x)(0.16)=(40)(0.15)
0.12x+6.4-0.16x=6
-0.04x+6.4=6
-0.04x=6-6.4
-0.04x=-0.4
-0.04x/-0.4=-0.4/-0.04
x=10

Amount = x = 10L
Concentration =12% =0.12
Solution B
Amount = 40-x =40-10=30L
Concentration =16% = 0.16
10L of 12% and 30L of 16% acid solutions were used to make 40L 15% Solution.